# Forecasting the Evolving Composition of Inbound Tourism Demand: A Bayesian Compositional Time Series Approach Using Platform Booking Data

### v1

@@ 5.3 Statistical Significance
Table 6 reports Diebold-Mariano test results comparing BDARMA forecasts against each benchmark for EMEA. BDARMA significantly outperforms naïve forecasts ($p=0.036$) and seasonal naïve ($p=0.027$). The improvement over ETS is marginally significant ($p=0.081$), while the comparison with SARIMA is not statistically significant ($p=0.329$).
@@
BDARMA vs. Naïve 1.84-1.84 0.036 Yes
BDARMA vs. SNaïve 1.99-1.99 0.027 Yes
BDARMA vs. ETS 1.43-1.43 0.081 No
BDARMA vs. SARIMA 0.45-0.45 0.329 No
Note: One-sided tests at $\alpha=0.05$. See Appendix A for implementation details.
Across all destination regions, BDARMA significantly outperforms seasonal naïve ($p<0.05$), reflecting the breakdown of annual seasonal patterns during the pandemic recovery period. For APAC, BDARMA shows marginally significant improvement over SARIMA ($p=0.075$), while comparisons with naïve and ETS are not significant.
@@ 5.4 Forecast Visualization
Figure 7 displays BDARMA forecasts against actuals for EMEA, focusing on the four largest origin markets (France, Great Britain, Germany, and the United States). The model captures the post-pandemic recovery dynamics, with forecast intervals (80% credible regions shown as shaded bands) providing principled uncertainty quantification derived from the Dirichlet predictive distribution.
Figure 7: BDARMA(1,1) forecasts versus actuals for EMEA. Solid lines show historical compositions; dashed lines show point forecasts; shaded regions indicate 80% prediction intervals; points show realized values. The model captures the post-pandemic stabilization of origin market shares.
@@ 5.5 Component-Level Accuracy
Table 7 reports MAE by origin market component for EMEA. BDARMA achieves the best accuracy for the three largest origin markets: France (28% improvement over naïve), Great Britain (51% improvement), and the United States (34% improvement). These substantial gains for major components drive BDARMA’s overall advantage.
@@
FR 0.0127 0.0178 0.0122 0.0148 0.0103 0.0092
GB 0.0136 0.0073 0.0075 0.0083 0.0098 0.0067
DE 0.0056 0.0063 0.0056 0.0049 0.0043 0.0053
US 0.0064 0.0150 0.0100 0.0069 0.0078 0.0042
ES 0.0025 0.0030 0.0025 0.0025 0.0029 0.0024
IT 0.0036 0.0030 0.0025 0.0040 0.0020 0.0032
NL 0.0030 0.0018 0.0016 0.0015 0.0021 0.0015
Other 0.0100 0.0142 0.0139 0.0097 0.0077 0.0091
Note: Bold indicates best performance for each component. BDARMA achieves the best MAE for four of eight components (FR, GB, US, ES, NL tied with ETS), with particularly strong gains for the three largest origin markets.
@@ 5.6 Pandemic Structural Break
Figure 8 illustrates the pandemic’s impact on origin market compositions by plotting deviations from pre-pandemic baseline levels (January 2019–February 2020 average) for the three most-affected origin markets in each destination region. The figure reveals substantial heterogeneity in how the pandemic reshaped tourism flows.
APAC experienced the most dramatic compositional shift: Chinese outbound travel collapsed by over 30 percentage points as China maintained strict travel restrictions, while South Korean and Australian origin shares increased to partially fill the gap. EMEA saw a surge in France-origin bookings of nearly 25 percentage points during the initial lockdown period, reflecting increased within-region travel when international borders closed. NAMER, already dominated by U.S.-origin bookings, saw this dominance intensify by an additional 15 percentage points as Canadian cross-border travel declined. LATAM exhibited a temporary spike in Brazil-origin bookings followed by gradual normalization.
Figure 8: Pandemic impact on origin market composition by destination region. Lines show changes in market share relative to the January 2019–February 2020 baseline for the three most-affected origin markets in each region. The vertical dashed red line indicates March 2020 (WHO pandemic declaration). Note the heterogeneous responses: APAC saw Chinese share collapse by 30 percentage points; EMEA experienced a France-origin surge of 25 percentage points; NAMER’s U.S.-origin dominance intensified; LATAM showed temporary Brazil-origin gains followed by normalization.
@@ 6.1 Summary of Findings
Our analysis demonstrates that BDARMA models achieve the lowest average forecast error across all destination regions, with particularly strong performance for destinations exhibiting meaningful compositional dynamics. For EMEA, BDARMA achieves 23% lower forecast error than naïve methods, with statistically significant improvement ($p=0.036$). The model also significantly outperforms seasonal naïve across all regions, reflecting its ability to adapt to regime changes while still capturing underlying temporal patterns.
@@
The BDARMA(1,1) specification consistently achieves the best in-sample fit across all destination regions, as measured by LOO-CV, indicating that both autoregressive and moving average components capture important features of compositional dynamics. The relative advantage of BDARMA varies systematically with the nature of compositional variation: it excels where multiple origin markets compete with shares in the 5–25% range (EMEA, APAC), performs comparably to simpler methods where one market dominates (NAMER), and is outperformed by ETS in settings with smooth trend dynamics (LATAM). As shown in Section 4, the HHI concentration analysis (Figure 2) helps explain this pattern: BDARMA provides greatest value for destinations with diverse origin portfolios where compositional dynamics are meaningful.
@@ 6.2 Implications for Destination Management
Seasonality management. The Fourier terms in our BDARMA specification capture regular seasonal patterns in both origin composition and compositional volatility—for example, increased European travel to Mediterranean destinations during summer months with tighter concentration around expected patterns. Understanding these patterns enables better alignment of marketing campaigns, staffing, and infrastructure utilization with anticipated demand composition.
Long-term strategic planning. While our forecast evaluation focused on 1–3 month horizons, the BDARMA framework can generate longer-term probabilistic projections. These can inform strategic decisions about market development, language capabilities, and partnership strategies with origin-market travel intermediaries.
@@ 6.3 Methodological Contributions
Second, we show that modeling seasonal variation in precision—not just mean composition—substantially improves forecast accuracy. This finding has implications beyond tourism, suggesting that compositional time series models in other domains may benefit from time-varying concentration parameters.
Third, we provide a systematic comparison of BDARMA against standard benchmarks using rolling origin evaluation. The finding that BDARMA achieves the lowest average error across all destination regions, and significantly outperforms transformed approaches (ETS and SARIMA on ILR data) for EMEA, demonstrates that the direct compositional modeling approach offers practical benefits beyond theoretical elegance.
@@ 7 Conclusion
This paper developed and applied Bayesian Dirichlet autoregressive moving average models to forecast the evolving composition of guest origin markets using Airbnb booking data. Our analysis of 96 months of reservations across four major destination regions demonstrates that BDARMA models achieve the lowest average forecast error across all destinations, with 23% lower error than naïve methods for EMEA and significant improvements over seasonal naïve benchmarks across all regions.
The COVID-19 pandemic caused dramatic compositional shifts—most notably the collapse of Chinese outbound travel and the surge in within-region bookings—with recovery trajectories that varied markedly across regions. Our models capture these dynamics through autoregressive and moving average terms while the Dirichlet likelihood ensures forecasts remain valid probability distributions. A key finding is that incorporating seasonal variation in the precision parameter substantially improves forecast accuracy.
@@ A.2 Diebold-Mariano Test Implementation
With $n=45$ forecast observations and $h=3$ step maximum horizon, we use bandwidth 1. Given the moderate sample size, we report $p$-values from the $t_{n-1}$ distribution rather than the asymptotic normal, following the finite-sample correction recommended by Diebold and Mariano (1995).

### v2

@@ 5.3 Statistical Significance
Table 6 reports Diebold-Mariano test results comparing BDARMA forecasts against each benchmark for EMEA. BDARMA significantly outperforms naïve forecasts ($p<0.001$), seasonal naïve ($p=0.013$), and ETS ($p<0.001$). The comparison with SARIMA is borderline ($p=0.058$), reflecting that both methods model temporal dynamics in the log-ratio space, with BDARMA’s advantage coming primarily from the Dirichlet likelihood and seasonal precision specification.
@@
BDARMA vs. Naïve 5.93-5.93 <0.001<0.001 Yes
BDARMA vs. SNaïve 2.26-2.26 0.013 Yes
BDARMA vs. ETS 7.34-7.34 <0.001<0.001 Yes
BDARMA vs. SARIMA 1.59-1.59 0.058 No
Note: One-sided tests at $\alpha=0.05$. Componentwise MAE loss with Newey–West HAC estimator (bandwidth = $\lfloor h^{1/3}\rfloor=1$). $n=84$ forecast observations, $df=83$. See A for implementation details.
Across destination regions, BDARMA significantly outperforms seasonal naïve for EMEA ($p=0.013$), NAMER ($p=0.001$), and APAC ($p<0.001$). For APAC, BDARMA significantly outperforms SARIMA ($p=0.003$) but not naïve or ETS. For LATAM, BDARMA significantly outperforms naïve ($p<0.001$) but not ETS or SARIMA, consistent with ETS achieving the lowest point accuracy for that region.
@@ 5.4 Forecast Visualization
Figure 7 displays an illustrative BDARMA forecast from the first evaluation origin (January 2022) for EMEA, showing the four largest named origin markets (France, Great Britain, Germany, and the United States) with 6-month-ahead point forecasts and 80% prediction intervals. The model captures the post-pandemic recovery dynamics, with forecast intervals (80% credible regions shown as shaded bands) providing principled uncertainty quantification derived from the Dirichlet predictive distribution.
Figure 7: Illustrative BDARMA(2,1) forecast from the first evaluation origin (January 2022) for EMEA, showing the four largest named origin markets (France, Great Britain, Germany, United States). Solid lines show historical compositions; dashed lines show 6-month-ahead point forecasts; shaded regions indicate 80% prediction intervals; points show realized values.
@@ 5.5 Component-Level Accuracy
Table 7 reports MAE by origin market component for EMEA. BDARMA achieves particularly strong accuracy for the United States component (44% improvement over naïve) and the Netherlands (50% improvement), while remaining competitive across other origins. The overall EMEA advantage is driven by consistent near-best performance across components rather than dominance on any single origin.
@@
FR 0.0110 0.0149 0.0128 0.0189 0.0095 0.0103
GB 0.0147 0.0065 0.0083 0.0092 0.0112 0.0075
DE 0.0067 0.0050 0.0057 0.0049 0.0049 0.0059
US 0.0087 0.0136 0.0101 0.0093 0.0096 0.0049
ES 0.0024 0.0028 0.0025 0.0033 0.0034 0.0036
IT 0.0039 0.0026 0.0026 0.0044 0.0021 0.0034
NL 0.0026 0.0018 0.0018 0.0016 0.0022 0.0013
Other 0.0157 0.0141 0.0151 0.0143 0.0091 0.0113
Note: Bold indicates best performance for each component. BDARMA achieves the best MAE for two of eight components (US, NL). SARIMA leads for FR, IT, and Other; SNaïve for GB; ETS/SARIMA tie for DE; and naïve for ES.
@@ 5.6 Pandemic Structural Break
Figure 8 illustrates the pandemic’s impact on origin market compositions by plotting deviations from pre-pandemic baseline levels (January 2019–February 2020 average) for the three most-affected origin components in each destination region. The figure reveals substantial heterogeneity in how the pandemic reshaped tourism flows.
APAC experienced the most dramatic compositional shift: Chinese outbound travel collapsed by over 30 percentage points as China maintained strict travel restrictions, while South Korean and other non-Chinese origin shares increased to partially fill the gap. EMEA saw a surge in France-origin bookings of nearly 25 percentage points during the initial lockdown period, reflecting increased within-region travel when international borders closed. NAMER, already dominated by U.S.-origin bookings, saw this dominance intensify by an additional 15 percentage points as Canadian cross-border travel declined. LATAM exhibited a temporary spike in Brazil-origin bookings followed by gradual normalization.
Figure 8: Pandemic impact on origin market composition by destination region. Lines show changes in market share relative to the January 2019–February 2020 baseline for the three most-affected origin components in each region. The vertical dashed red line indicates March 2020 (WHO pandemic declaration). Note the heterogeneous responses: APAC saw Chinese share collapse by 30 percentage points; EMEA experienced a France-origin surge of 25 percentage points; NAMER’s U.S.-origin dominance intensified; LATAM showed temporary Brazil-origin gains followed by normalization.
@@ 6.1 Summary of Findings
Our analysis demonstrates that BDARMA models achieve the lowest forecast error for EMEA and competitive performance across other destination regions, with particular strength where multiple origin markets exhibit rich compositional dynamics. For EMEA, BDARMA achieves 27% lower forecast error than naïve methods, with highly significant improvement ($p<0.001$). Averaged across all destinations, ETS slightly edges BDARMA (MAE 0.0086 vs. 0.0089), though the difference is not statistically significant. The model also significantly outperforms seasonal naïve across most regions, reflecting its ability to adapt to regime changes while still capturing underlying temporal patterns.
@@
The BDARMA(1,1) specification achieves the best in-sample fit for three of four destination regions (BDARMA(2,1) is preferred for EMEA), as measured by LOO-CV, indicating that both autoregressive and moving average components capture important features of compositional dynamics. The relative advantage of BDARMA varies systematically with the nature of compositional variation: it excels where multiple origin markets compete with shares in the 5–25% range (EMEA), performs comparably to simpler methods where one market dominates (NAMER), and is outperformed by ETS in settings with smooth trend dynamics (LATAM). As shown in Section 4, the HHI concentration analysis (Figure 2) helps explain this pattern: BDARMA provides greatest value for destinations with diverse origin portfolios where compositional dynamics are meaningful.
@@ 6.2 Implications for Destination Management
Seasonality management. The Fourier terms in our BDARMA specification capture regular seasonal patterns in both origin composition and compositional volatility, for example, increased European travel to Mediterranean destinations during summer months with tighter concentration around expected patterns. Understanding these patterns enables better alignment of marketing campaigns, staffing, and infrastructure utilization with anticipated demand composition.
Long-term strategic planning. While our forecast evaluation focused on 1–6 month horizons, the BDARMA framework can generate longer-term probabilistic projections. These can inform strategic decisions about market development, language capabilities, and partnership strategies with origin-market travel intermediaries.
@@ 6.3 Methodological Contributions
Second, we show that modeling seasonal variation in precision, not just mean composition, substantially improves forecast accuracy. This finding has implications beyond tourism, suggesting that compositional time series models in other domains may benefit from time-varying concentration parameters.
Third, we provide a systematic comparison of BDARMA against standard benchmarks using rolling origin evaluation. The finding that BDARMA achieves the lowest error for EMEA and significantly outperforms naïve, seasonal naïve, and ETS ($p<0.001$, $p=0.013$, and $p<0.001$ respectively), demonstrates that the direct compositional modeling approach offers practical benefits beyond theoretical elegance, particularly for destinations with diverse origin portfolios.
@@ 7 Conclusion
This paper developed and applied Bayesian Dirichlet autoregressive moving average models to forecast the evolving composition of guest origin markets using Airbnb booking data. Our analysis of 108 months of reservations across four major destination regions demonstrates that BDARMA models achieve 27% lower forecast error than naïve methods for EMEA ($p<0.001$) and significant improvements over seasonal naïve benchmarks across most regions. Averaged across destinations, BDARMA performs comparably to the best-performing benchmark (ETS), with its advantage concentrated in compositionally diverse markets.
The COVID-19 pandemic caused dramatic compositional shifts, most notably the collapse of Chinese outbound travel and the surge in within-region bookings, with recovery trajectories that varied markedly across regions. Our models capture these dynamics through autoregressive and moving average terms while the Dirichlet likelihood ensures forecasts remain valid probability distributions. A key finding is that incorporating seasonal variation in the precision parameter substantially improves forecast accuracy.
@@ A.2 Diebold-Mariano Test Implementation
With $n=84$ forecast observations (14 origins $\times$ 6 horizons) and $h=6$ step maximum horizon, we use bandwidth $\lfloor 6^{1/3}\rfloor=1$. Given the moderate sample size, we report $p$-values from the $t_{n-1}$ distribution rather than the asymptotic normal, following the finite-sample correction recommended by Diebold and Mariano (1995).